# Uncertainty Quantification and Confidence Calibration in Large Language Models: A Survey

### v1

@@ Abstract.
Large Language Models (LLMs) excel in text generation, reasoning, and decision-making, enabling their adoption in high-stakes domains such as healthcare, law, and transportation. However, their reliability is a major concern, as they often produce plausible but incorrect responses. Uncertainty quantification (UQ) enhances trustworthiness by estimating confidence in outputs, enabling risk mitigation and selective prediction. However, traditional UQ methods struggle with LLMs due to computational constraints and decoding inconsistencies. Moreover, LLMs introduce unique uncertainty sources, such as input ambiguity, reasoning path divergence, and decoding stochasticity, that extend beyond classical aleatoric and epistemic uncertainty. To address this, we introduce a new taxonomy that categorizes UQ methods based on computational efficiency and uncertainty dimensions (input, reasoning, parameter, and prediction uncertainty). We evaluate existing techniques, assess their real-world applicability, and identify open challenges, emphasizing the need for scalable, interpretable, and robust UQ approaches to enhance LLM reliability.
@@
copyright: none
@@ 1. Introduction
Large Language Models (LLMs) like GPT-4 (Achiam et al., 2023) and PaLM (Anil et al., 2023) have achieved remarkable capabilities in text generation, reasoning, and decision-making, driving their adoption in high-stakes domains such as healthcare diagnostics (Qiu et al., 2024; Ullah et al., 2024; Goh et al., 2024; Da et al., 2024d), legal analysis (Cheong et al., 2024; Jayakumar et al., 2023; Drápal et al., 2023), and transportation systems (Jin et al., 2023; Da et al., 2024b; Zheng et al., 2023a; Lai et al., 2023). However, their reliability remains a critical concern: LLMs often produce plausible but incorrect or inconsistent outputs, with studies showing that over 30% of answers in medical QA tasks contain factual errors (Jin et al., 2021). In sensitive applications, these limitations pose risks ranging from misinformation to life-threatening misdiagnoses, underscoring the urgent need for robust reliability frameworks.
@@
Uncertainty quantification (UQ) emerges as a pivotal mechanism to enhance LLM trustworthiness by explicitly modeling confidence in model outputs. By estimating uncertainty, users can identify low-confidence predictions for human verification, prioritize high-certainty responses, and mitigate risks like overconfidence in hallucinations (Slobodkin et al., 2023; Liu et al., 2024a). For instance, in clinical settings, uncertainty-aware LLMs could flag uncertain diagnoses for specialist review, reducing diagnostic errors by up to 41% (Sen et al., 2024). This capability is particularly critical as LLMs transition from experimental tools to production systems requiring accountability.
@@
Traditional UQ methods face significant hurdles when applied to Large Language Models (LLMs). Bayesian approaches like Monte Carlo dropout (Gal and Ghahramani, 2016) are computationally prohibitive for trillion-parameter models and natural language generation (NLG) tasks, while ensemble methods struggle with consistency across diverse decoding strategies (Lu et al., 2024). Furthermore, LLMs introduce unique uncertainty sources—such as input ambiguity, reasoning path divergence, and decoding stochasticity—that transcend classical aleatoric and epistemic categorizations (Hüllermeier and Waegeman, 2021). The complexity of LLMs, characterized by sequence generation over vast parameter spaces and reliance on massive datasets, exacerbates uncertainty challenges. This complexity, coupled with the critical need for reliable outputs in high-stakes applications, positions UQ for LLMs as a compelling yet underexplored research frontier.
@@
This tutorial introduces a novel taxonomy for LLM uncertainty quantification, categorizing methods along two axes: (1) computational efficiency (e.g., single-pass vs. sampling-based techniques) and (2) uncertainty dimensions (input, reasoning, parametric, predictive). Our framework addresses three gaps in prior work: First, it decouples uncertainty sources unique to LLMs from traditional ML contexts. Second, it evaluates methods through the lens of different dimensions of the responses from LLM: input uncertainty, reasoning uncertainty, parameter uncertainty, and prediction uncertainty. Each of these dimensions may involve aleatoric uncertainty, epistemic uncertainty, or a mixture of both. Third, it identifies understudied areas like reasoning uncertainty, which accounts for 58% of errors in multi-step QA tasks (Da et al., 2025a).
Contributions: We provide (1) the first systematic review of UQ methods tailored to different dimensions of uncertainty in LLM, including input, reasoning, parameter, and prediction uncertainty; (2) a comprehensive introduction to the evaluation of UQ methods and the challenges.
@@
Connection to Existing Surveys: Prior surveys (Shorinwa et al., 2024; Huang et al., 2024b; Beigi et al., 2024; Huang et al., 2025a; Rawte et al., 2023; Zhang et al., 2023) focus narrowly on hallucination detection or retrofitting classical UQ taxonomies, neglecting LLM-specific challenges like prompt-driven input uncertainty. Our work uniquely addresses the interplay between model scale, open-ended generation, and uncertainty dynamics—factors critical for modern LLMs but overlooked in earlier frameworks.
@@
The remainder of this survey is structured as follows: Section 2 characterizes LLM uncertainty dimensions and differentiates confidence from uncertainty. Section 3 evaluates UQ methods using our taxonomy. Section 4 introduces the evaluation of UQ methods for LLM. Sections 5 and 6 introduce the applications of UQ in different domains with LLMs and identify open challenges and future directions. Table 1 shows the notation and Table 2 shows an overview of all methods we discuss in this paper.
@@ 2.1. Sources of Uncertainty in LLMs
For uncertainty quantification of traditional machine learning tasks such as classification or regression (Young et al., 2024), there are mainly two types of uncertainty (Der Kiureghian and Ditlevsen, 2009): aleatoric and epistemic uncertainty. Aleatoric uncertainty defines the uncertainty from noise in the dataset. Epistemic uncertainty, on the other hand, arises from the model’s lack of knowledge about the underlying data distribution.
Aleatoric uncertainty in LLMs primarily stems from data sources used to train LLMs, which contain inconsistencies, biases, and contradicting information. Furthermore, ambiguity in natural language contributes to aleatoric uncertainty, as different interpretations of the same prompt can lead to multiple plausible responses.
On the other hand, When encountering unfamiliar topics, LLMs may exhibit high epistemic uncertainty, often manifesting as hallucinations or overconfident yet incorrect statements. Epistemic uncertainty can be reduced through domain-specific fine-tuning or retrieval-augmented generation techniques that allow the model to access external knowledge sources.
Though uncertainty for LLMs can also be classified as these two categories, these two categories alone are insufficient to fully capture the complexities of uncertainty in the language model. The unique nature of LLM inference introduces additional uncertainty factors. In particular, LLMs exhibit uncertainty not only due to training data limitations but also due to input variability and decoding mechanisms. Therefore, to address these challenges, we introduce four new dimensions of uncertainty.
@@ 2.1.2. Uncertainty with Different Dimensions
To systematically analyze uncertainty in LLMs, we categorize it into four key dimensions: input uncertainty, reasoning uncertainty, parameter uncertainty, and prediction uncertainty. Each dimension may involve aleatoric uncertainty, epistemic uncertainty, or a combination of both. This structured framework provides a more comprehensive understanding of uncertainty quantification in LLMs.
@@
Input Uncertainty (Aleatoric Uncertainty): Input uncertainty arises when a prompt is ambiguous or underspecified, making it impossible for an LLM to generate a single definitive response. This is inherently aleatoric, as even a “perfect model” cannot resolve the ambiguity. For instance, “What is the capital of this country?” lacks sufficient context, leading to unpredictable outputs. Similarly, “Summarize this document” may yield different responses depending on different expected details.
@@
Reasoning Uncertainty (Mixed Uncertainty): Reasoning uncertainty occurs when an LLM derives answers through multi-step logic or retrieval, leading to ambiguous or incorrect reasoning. This uncertainty is aleatoric when the problem itself is ambiguous and epistemic when the model cannot offer robust reasoning.
@@
Parameter Uncertainty (Epistemic Uncertainty): Parameter uncertainty stems from gaps in the training data, where the model has either never seen relevant information or has learned an incorrect representation. Unlike aleatoric uncertainty, epistemic uncertainty can be reduced by improving the model’s knowledge base. Bayesian methods (Gal and Ghahramani, 2016), deep ensembles (Lakshminarayanan et al., 2017), and uncertainty-aware training (Mukherjee and Awadallah, 2020) can help quantify and mitigate this type of uncertainty.
@@
Prediction Uncertainty(Mixed Uncertainty): Prediction uncertainty refers to variability in generated outputs across different sampling runs, influenced by both aleatoric and epistemic sources. For example, when asked, “What are the side effects of a new experimental drug?” the model’s responses might vary significantly across different sampling runs, especially if no reliable data is available in its training set. A high-variance output distribution in such scenarios suggests that the model is both aware of multiple possible answers, reflecting aleatoric uncertainty, and uncertain due to incomplete knowledge, highlighting epistemic uncertainty.
@@
$x$ The question that LLMs answer
$s$ Generation from LLMs
$w_{i}$ i-token in the generation $s$
$\mathcal{D}$ Dictionary of LLMs
$U(x)$ Uncertainty of question $x$
$C(x,s)$ Confidence of generation $s$ given $x$
$H(s)$ Entropy of generation $s$
Table 1. Notation used in this paper.
2.2. Uncertainty and Confidence in LLMs
Uncertainty quantification and confidence estimation are closely related yet distinct concepts in the context of large language models (LLMs). Uncertainty is a property of the model’s predictive distribution, capturing the degree of variability or unpredictability given a particular input. In contrast, confidence reflects the model’s belief in the correctness of a particular answer or prediction. As a concrete example, in the context of classification, a simple confidence measure is the predicted probability $\hat{p}(Y=y|x)$ (an uncertainty measure which does not depend on the particular prediction $y$ could be entropy, taking the form of $\sum_{y}-\hat{p}(Y=y|x)\log{\hat{p}(Y=y|x)}$). The corresponding confidence score in NLG (for an auto-regressive LM) is the joint probability for the generated sequence:
The log of Eq. 1 is sometimes referred to as sequence likelihood (Zhao et al., 2023a). In general, while an uncertainty estimate takes the form of $U(x)$, confidence estimates could be expressed as $C(\mathbf{x},\mathbf{s})$. Note that unlike classification tasks, not all NLG applications have the notion of a “correct” answer (e.g. summarization). Thus, while for the ease of writing we use the term correctness throughout this section, it should really be interpreted as the gold-label for the particular application. Note also that in most cases, the correct answer is not unique, and thus such gold-label typically takes the form of a “correctness function” that decides whether a particular generation $\mathbf{s}$ is good or not. We will denote such a function as $f(\mathbf{s}|\mathbf{x})$.
@@ 2.2. Uncertainty and Confidence in LLMs
There are usually two dimensions along which researchers improve confidence estimates in NLG, which is unsurprisingly largely influenced by confidence scoring literature from classification (Jiang et al., 2018; Corbière et al., 2019), especially binary classification. We refer to them as ranking performance and calibration:
@@
  • Ranking performance refers to the discriminative power of the confidence measure on the correctness. Like in classification, LLM confidence is often evaluated by its ability to separate correct and incorrect answers, thus typically measured by evaluation metrics like AUROC (Kadavath et al., 2022) or AUARC (Lin et al., 2024) as detailed in Section 4.
  • Calibration refers to closing the gap between the confidence score and the expected correctness conditioned on confidence score. It has a long history preceding LLM or even modern machine learning (Murphy and Winkler, 1977; Degroot and Fienberg, 1983; Murphy and Winkler, 1984), but bears slightly different meanings in NLP. In general, we could define a perfectly calibrated confidence measure to achieve the following:
    (2) $\forall c,\mathbb{E}[f(\mathbf{s}|\mathbf{x})|C(\mathbf{x},\mathbf{s})=c]=c,$
@@
    where the expectation is taken over the joint distribution of $\mathbf{x}$ and generation $\mathbf{s}$. A lot of papers focus on evaluating the calibration quality of specific LMs and tasks (Wang et al., 2020; Khojah et al., 2022; Kumar and Sarawagi, 2019). Evaluation typically relies on variants of Expected Calibration Error (ECE) (Tian et al., 2023; Kumar and Sarawagi, 2019; Li et al., 2024b). Oftentimes confidence scores from classification could be directly applied (Stengel-Eskin and Van Durme, 2023; Zhao et al., 2021; Jiang et al., 2021) in order to evaluate whether an LM is over- or under-confident, especially for de facto classification tasks like sentiment analysis or multiple-choice QA.

### v2

@@ Abstract.
Uncertainty quantification (UQ) enhances the reliability of Large Language Models (LLMs) by estimating confidence in outputs, enabling risk mitigation and selective prediction. However, traditional UQ methods struggle with LLMs due to computational constraints and decoding inconsistencies. Moreover, LLMs introduce unique uncertainty sources, such as input ambiguity, reasoning path divergence, and decoding stochasticity, that extend beyond classical aleatoric and epistemic uncertainty. To address this, we introduce a new taxonomy that categorizes UQ methods based on computational efficiency and uncertainty dimensions, including input, reasoning, parameter, and prediction uncertainty. We evaluate existing techniques, summarize existing benchmarks and metrics for UQ, assess their real-world applicability, and identify open challenges, emphasizing the need for scalable, interpretable, and robust UQ approaches to enhance LLM reliability.
@@
copyright: acmlicensedjournalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3736569isbn: 979-8-4007-1454-2/2025/08ccs: Computing methodologies Machine learningccs: Computing methodologies Natural language processing
@@ 1. Introduction
Large Language Models (LLMs) like GPT-4 (Achiam et al., 2023) have achieved remarkable capabilities in text generation, reasoning, and decision-making, driving their adoption in high-stakes domains such as healthcare diagnostics (Qiu et al., 2024; Da et al., 2024d), legal analysis (Cheong et al., 2024; Li et al., 2024b), and transportation systems (Da et al., 2024b; Lai et al., 2025; Xing et al., 2025). However, their reliability remains a critical concern: LLMs often produce plausible but incorrect or inconsistent outputs, with studies showing that over 30% of answers in medical QA tasks contain factual errors (Jin et al., 2021). In sensitive applications, these limitations pose risks ranging from misinformation to life-threatening misdiagnoses, underscoring the urgent need for robust reliability frameworks.
@@
Uncertainty quantification (UQ) emerges as an important mechanism to enhance LLM reliability by explicitly modeling confidence in model outputs. By estimating uncertainty, users can identify low-confidence predictions for human verification, prioritize high-certainty responses, and mitigate risks like overconfidence in hallucinations (Liu et al., 2024a). For instance, in clinical settings, uncertainty-aware LLMs could flag uncertain diagnoses for specialist review, reducing diagnostic errors by up to 41% (Sen et al., 2024). This capability is particularly critical as LLMs’ transition from experimental tools to production systems requiring accountability.
@@
Traditional UQ methods face significant hurdles when applied to LLMs. Bayesian approaches like Monte Carlo dropout (Gal and Ghahramani, 2016) are computationally prohibitive for trillion-parameter models and natural language generation (NLG) tasks, while ensemble methods struggle with consistency across diverse decoding strategies (Lu et al., 2024). Furthermore, LLMs introduce unique uncertainty sources, such as input ambiguity (Chen et al., 2025a; Guo et al., 2021), reasoning path divergence, and decoding stochasticity that transcend classical aleatoric and epistemic categorizations (Hüllermeier and Waegeman, 2021). The complexity of LLMs, characterized by sequence generation over vast parameter spaces and reliance on massive datasets, exacerbates uncertainty challenges. This complexity, coupled with the critical need for reliable outputs in high-stakes applications, positions UQ for LLMs as a compelling yet underexplored research frontier.
@@
Targeting the unique challenges of UQ in LLMs, this survey firstly introduces a novel taxonomy for LLM UQ, categorizing methods along two axes: (1) computational efficiency (e.g., single-pass vs. sampling-based techniques) and (2) uncertainty dimensions (input, reasoning, parametric, predictive). This framework addresses three gaps in prior works: First, it decouples uncertainty sources unique to LLMs from traditional ML contexts. Second, it evaluates methods through the lens of different dimensions of the responses from LLM: input uncertainty, reasoning uncertainty, parameter uncertainty, and prediction uncertainty. Each of these dimensions may involve aleatoric uncertainty, epistemic uncertainty, or a mixture of both. Third, it identifies understudied areas like reasoning uncertainty, challenges, and possible future directions.
@@
Connection to Existing Surveys: Prior surveys (Shorinwa et al., 2024; Huang et al., 2024b, 2025a) focus on hallucination detection or retrofitting classical UQ taxonomies, neglecting LLM-specific challenges like prompt-driven input uncertainty. Our work uniquely addresses the interplay between model scale, open-ended generation, and uncertainty dynamics, which are critical for modern LLMs but overlooked in earlier frameworks.
@@
The remainder of this survey is structured as follows: Section 2 characterizes LLM uncertainty dimensions and differentiates confidence from uncertainty. Section 3 evaluates UQ methods using our taxonomy. Section 4 introduces the evaluation of UQ methods for LLM, including benchmarks and metrics. Sections 5 and 6 introduce the applications of UQ in different domains with LLMs and identify open challenges and future directions.
@@ 2.1. Sources of Uncertainty in LLMs
For UQ on traditional machine learning tasks such as classification or regression (Young et al., 2025), there are mainly two types of uncertainty (Der Kiureghian and Ditlevsen, 2009; Ye et al., 2024): aleatoric uncertainty, which models the uncertainty from noise in the dataset, and epistemic uncertainty, which arises from the model’s lack of knowledge about the underlying data distribution.
Aleatoric uncertainty in LLMs primarily stems from data sources used to train LLMs, which contain inconsistencies, biases, and contradicting information. Additionally, ambiguity in natural language also contributes to aleatoric uncertainty, as different interpretations of the same prompt can lead to multiple plausible responses. On the other hand, when encountering unfamiliar topics, LLMs may exhibit high epistemic uncertainty, often manifesting as hallucinations or overconfident yet incorrect statements. Epistemic uncertainty can be reduced through domain-specific fine-tuning or retrieval-augmented generation techniques that allow the model to access external knowledge sources.
@@ 2.1.2. Uncertainty with Different Dimensions
While the uncertainty for LLMs can also be classified through aleatoric and epistemic uncertainty, these two categories alone are insufficient to fully capture the complexities of uncertainty in LLMs. In particular, LLMs exhibit uncertainty not only due to training data limitations but also due to input variability and decoding mechanisms. Therefore, in the following, we formulate four dimensions of uncertainty, each of which may involve aleatoric uncertainty, epistemic uncertainty, or a combination of both. The technical methods on how to quantify these uncertainties will be discussed later in Section 3.
@@
$\bullet$ Input Uncertainty (Aleatoric Uncertainty): Input uncertainty arises when a prompt is ambiguous or underspecified, making it impossible for an LLM to generate a single definitive response. This is inherently aleatoric, as even a “perfect model” cannot resolve the ambiguity. For instance, “What is the capital of this country?” lacks sufficient context, leading to unpredictable outputs. Similarly, “Summarize this document” may yield different responses depending on different expected details.
@@
$\bullet$ Reasoning Uncertainty (Mixed Uncertainty): Reasoning uncertainty occurs when an LLM derives answers through multi-step reasoning (Mondorf and Plank, 2024) or retrieval (Li et al., 2024c), where the uncertainty of each step can lead to ambiguous or incorrect results. This uncertainty is aleatoric when the problem itself is ambiguous and epistemic when the model cannot offer robust reasoning.
@@
$\bullet$ Parameter Uncertainty (Epistemic Uncertainty): Parameter uncertainty stems from training data gaps, where the model has either never seen relevant information or has learned an incorrect representation. Unlike aleatoric uncertainty, epistemic uncertainty can be reduced by improving the model’s knowledge base. Bayesian methods (Gal and Ghahramani, 2016), deep ensembles (Lakshminarayanan et al., 2017), and uncertainty-aware training (Mukherjee and Awadallah, 2020) can help quantify and mitigate this type of uncertainty.
@@
$\bullet$ Prediction Uncertainty (Mixed Uncertainty): Prediction uncertainty refers to variability in generated outputs across different sampling runs, influenced by both aleatoric and epistemic sources. For example, when asked “What are the side effects of a new experimental drug?”, the model’s responses might vary significantly across different sampling runs, especially if no reliable data is available in its training set. A high-variance output distribution in such scenarios suggests that the model is both aware of multiple possible answers, reflecting aleatoric uncertainty, and uncertain due to incomplete knowledge, highlighting epistemic uncertainty.
2.2. Uncertainty and Confidence in LLMs
2.2.1. Classical Confidence Estimation
UQ and confidence estimation are closely related yet distinct concepts. In traditional machine learning, uncertainty is a property of the model’s predictive distribution, capturing the degree of variability or unpredictability given a particular input. In contrast, confidence reflects the model’s belief in the correctness of a particular prediction. If we follow the definition in classification tasks, the confidence measure would be the predicted probability $\hat{p}(Y=y|x)$ given input $x$ (an uncertainty measure which does not depend on the particular prediction $y$ could be entropy, taking the form of $\sum_{y}-\hat{p}(Y=y|x)\log{\hat{p}(Y=y|x)}$). Table 1 shows a similar notation in a question-answering (QA) task in Natural Language Generation (NLG). The corresponding confidence score in NLG tasks for an auto-regressive language model would be the joint probability for the generated sequence:
The log of Eq. 1 is sometimes referred to as sequence likelihood (Zhao et al., 2023a). In general, an uncertainty estimate in existing literature usually takes the form of $U(\mathbf{x})$, while confidence estimates are usually expressed as $C(\mathbf{x},\mathbf{s})$. Note that, unlike classification tasks, not all NLG applications have the notion of a “correct” answer (e.g., summarization). Thus, while for the ease of writing we use the term correctness throughout this section, it should really be interpreted as the gold-label for the particular application. Note also that in most cases, the correct answer is not unique, and thus such gold-label typically takes the form of a “correctness function” that decides whether a particular generation $\mathbf{s}$ is good or not. We will denote such a function as $f(\mathbf{s}|\mathbf{x})$.
@@ 2.2.1. Classical Confidence Estimation
$\mathbf{x}$ The question that LLMs answer
$\mathbf{s}$ Generation from LLMs
$w_{i}$ i-th token in the generation $\mathbf{s}$
$\mathcal{D}$ Corpus of LLMs
$U(\mathbf{x})$ Uncertainty of question $\mathbf{x}$
$C(\mathbf{x},\mathbf{s})$ Confidence of generation $\mathbf{s}$ given $\mathbf{x}$
$H(\mathbf{s})$ Entropy of generation $\mathbf{s}$
Table 1. Notations used in this paper for an exemplary QA task.
2.2.2. Confidence Improvement
@@
There are usually two dimensions along which researchers improve confidence estimates in NLG, which is unsurprisingly largely influenced by confidence scoring literature from classification (Jiang et al., 2018), especially binary classification. We refer to them as ranking performance and calibration:
@@
$\bullet$ Ranking performance refers to the discriminative power of the confidence measure on the correctness. Like in classification, LLM confidence is often evaluated by its ability to separate correct and incorrect answers, thus typically measured by evaluation metrics like AUROC (Kadavath et al., 2022) or AUARC (Lin et al., 2024) as detailed in Section 4.
@@
$\bullet$ Calibration refers to closing the gap between the confidence score and the expected correctness conditioned on confidence score. It has a long history preceding even modern machine learning (Murphy and Winkler, 1977), but bears slightly different meanings in NLP. In general, we could define a perfectly calibrated confidence measure to achieve: $\forall c,\mathbb{E}[f(\mathbf{s}|\mathbf{x})|C(\mathbf{x},\mathbf{s})=c]=c,$ where the expectation is taken over the joint distribution of $\mathbf{x}$ and generation $\mathbf{s}$. A lot of papers focus on evaluating the calibration quality of specific language models (LMs) and tasks (Wang et al., 2020; Kumar and Sarawagi, 2019). Evaluation typically relies on variants of Expected Calibration Error (ECE) (Tian et al., 2023; Kumar and Sarawagi, 2019). Oftentimes confidence scores from classification could be directly applied (Stengel-Eskin and Van Durme, 2023) in order to evaluate whether an LM is over- or under-confident, especially for de facto classification tasks like sentiment analysis or multiple-choice QA.
2.2.3. Confidence Estimation for LLMs
Confidence estimation in large language models (LLMs) refers to the task of quantifying how certain a model is about a specific generated output. In this subsection, we review three major families of approaches to confidence estimation in LLMs: